fourteen is three times the difference of a number and 6. write an equations
Answer by anjali_yamini(27)

14 is 3 times the difference of anumber and 6.
let
the number = x
so, 3 * ( 6 - x) = 14
so, 6 - x = 14 / 3
so, 6 - 14/3 = x
so, (18 - 14) / 3 = x
so, 4/3= 1.3 =x

check :
3*(6 - 4/3)
= 3 * ( (18-4)/3)
= 3 * 14 /3
= 14